OpenStudy (anonymous):

Integral help y=cosx, y=2-cosx, 0

OpenStudy (anonymous):

There are two y's. you need integral of what??

OpenStudy (anonymous):

I'm trying to find the area in-between the 2 curves

OpenStudy (anonymous):

in that interval of 0 to 2pi

OpenStudy (anonymous):

Sorry I have no clue..

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

Can you show me your steps?? May be I can help you somewhat..

OpenStudy (anonymous):

\[2-\cos(x)\geq \cos(x)\] so your area between the curves is found via \[\int_0^{2\pi}2-\cos(x)-\cos(x)dx=\int_0^{2\pi}2-2\cos(x)\]

OpenStudy (anonymous):

should be straightforward to compute, it is \[\int_0^{2\pi}2dx-2\int_0^{2\pi}\cos(x)dx\] first part is \(2\pi\times 2=4\pi\) from your eyeballs, second one is what you get when you take the anti derivative of sine, and evaluate

OpenStudy (anonymous):

satellite73 is correct.

hero (hero):

By the way @satellite73, I usually use my brain and a pencil to evaluate things. Not necessarily eyeballs

OpenStudy (anonymous):

sure but not \(\int_a^b cdx=c(b-a)\)